# Non-normal Dynamics on Non-reciprocal Networks: Reactivity and Effective Dimensionality in Neural Circuits

### v1

@@ Abstract
Non-reciprocal interactions are a defining feature of many complex systems, biological, ecological, and technological, often pushing them far from equilibrium and enabling rich dynamical responses. These asymmetries can arise at multiple levels: locally, in the dynamics of individual units, and globally, in the topology of their interactions. In this work, we investigate how these two forms of non-reciprocity interact in networks of neuronal populations. At the local level, each population is modeled by a non-reciprocally coupled set of excitatory and inhibitory neural populations exhibiting transient amplification and reactivity. At the network level, these populations are coupled via directed, asymmetric connections that introduce structural non-normality. Since non-reciprocal interactions generically lead to non-normal linear operators, we frame both local and global asymmetries in terms of non-normal dynamics. Using a modified Wilson–Cowan framework, we analyze how the interplay between these two types of non-normality shapes the system’s behavior. We show that their combination leads to emergent collective dynamics, including fluctuation-driven transitions, dimensionality reduction, and novel nonequilibrium steady states. Our results provide a minimal yet flexible framework to understand how multi-scale non-reciprocities govern complex dynamics in neural and other interconnected systems.
@@ I Introduction
In this paper, we aim to bring together and clarify two distinct, but often coexisting, manifestations of non-reciprocity in complex systems, both of which can be naturally interpreted through the lens of non-normality: a mathematical property of matrices where eigenvectors are non-orthogonal, allowing for transient amplification of perturbations. While non-normality provides a technical language to describe these effects, we emphasize that it is the underlying physical non-reciprocity — whether in the local dynamics or the network architecture — that drives the phenomena of interest.
@@
The first form we address is dynamical non-normality, which arises at the level of the internal dynamics of individual units or subsystems. When the local Jacobian or stability matrix is non-normal, even stable fixed points can exhibit large transient responses to small perturbations, especially in the presence of noise [12, 13]. In neural circuits, for instance, populations of excitatory and inhibitory neurons often operate near critical balance, leading to strongly non-normal dynamics that transiently amplify fluctuations in input, contributing to fast signal processing, selectivity, and flexible population coding [14, 15, 16]. These effects are intrinsically linked to asymmetries in the internal coupling and can be viewed as a form of microscopic non-reciprocity between excitatory and inhibitory neurons.
@@
The second form is what we term structural non-normality, which stems from the network-level asymmetry in the way local populations are connected at a larger scale. In many real-world systems, especially biological and engineered networks, the connectivity matrix that governs inter-unit influences is inherently non-reciprocal. Neural circuits, for example, are not fully bidirectional: higher-order cortical areas send feedback to early sensory regions, but not necessarily with equal weight or latency [17, 18, 19]. Similarly, ecological and social systems often exhibit trophic or influence hierarchies, where some magnitude flows in preferred directions [20]. When such asymmetries are strong, the overall network becomes structurally non-normal, leading to directional instabilities, long-range correlations, waves, and flow-like dynamics.
The central hypothesis we aim to explore is whether the interplay between local dynamical non-normality and global structural non-normality gives rise to qualitatively new collective behaviors, beyond what either form could produce in isolation. In other words, when locally non-reciprocal units (e.g., excitatory/inhibitory local neural populations) are embedded in a globally non-reciprocal network (e.g., feedforward/feedback cortical circuits), the system exhibits emergent properties that are not captured by standard linear response theory or symmetric interaction models. These might include enhanced stochastic amplification, novel transient modes, and selective propagation of fluctuations and signals across the network.
@@
As a test of this hypothesis, here we analyze a minimal yet general neural model consisting of multiple interacting dynamical units, each described by a non-normal local system, coupled through a network, that in general is non-reciprocal. Using tools from linear stability analysis, stochastic dynamics, and spectral theory, we explore how the composition of non-reciprocities across scales governs the system’s susceptibility to noise, its capacity to transmit signals, and its repertoire of emergent dynamical states.
@@
While our focus is primarily on neural systems —where both forms of non-reciprocity are particularly notorious— the conceptual framework we propose is broadly applicable to other domains, including active materials, biological networks, and collective decision-making systems. In all these contexts, non-reciprocity serves as a dynamical and structural engine for complexity, and understanding its layered manifestations is key to unlocking the behavior of high-dimensional, complex systems.
A theoretical neuroscience perspective:
@@ II Wilson-Cowan model for neural populations
Dynamical non-normal effects at a single region
@@
[…]
where $x(t)$ and $y(t)$ are the activation levels of the excitatory and inhibitory neuronal populations, respectively. The parameter $\alpha$ represents the rate of spontaneous activity decay, while $s$ denotes the net incoming current, $s(t)=\gamma_{\mu}x(t)-\gamma_{\nu}y(t)+h$, which is simply the sum of all synaptic excitatory and inhibitory inputs, weighted by their respective synaptic strengths —-$(\gamma_{\mu}$ for excitation and $\gamma_{\nu}$ for inhibition— plus a constant external input current, $h$. To reduce the dimensionality of the parameter space, all synaptic weights of a given type have been set to the same value. Finally $f(s)$ is a sigmoid non-negative response function. The system displays two distinct dynamical regimes: a quiescent (inactive) phase, where $\bar{x}=\bar{y}=0$, and an active phase $\bar{x}=\bar{y}>0$, characterized by nonzero steady-state activity with excitation dominating the response ($\bar{x},\bar{y}\neq 0$). These phases are separated by a bifurcation that occurs at the critical point $\gamma_{\mu}=\gamma_{\nu}+\alpha$: below this threshold, the system remains in the quiescent state, while above it, activity becomes self-sustained (see below).
In the deterministic limit, using variables $\Sigma=(x+y)/2$ and $\Delta=(x-y)/2$, linearization around the active-state fixed point $(\bar{\Sigma},\bar{\Delta}\neq 0)$ yields a non-normal Jacobian matrix:
@@
where the eigenvalues are given by $\lambda_{1}=\alpha+f(\bar{s})+(1-\bar{\Sigma})\omega_{0}f^{\prime}(\bar{s})$ and $\lambda_{2}=\alpha+f(\bar{s})$, with $\bar{s}=\omega_{0}\bar{\Sigma}+h$. The feedforward term $\omega_{\text{ff}}=(1-\bar{\Sigma})(\gamma_{\mu}+\gamma_{\nu})f^{\prime}(\bar{s})$ governs the asymmetric influence of $\Sigma$ on $\Delta$, while the parameter $\omega_{0}=\gamma_{\mu}-\gamma_{\nu}$ quantifies the difference between excitatory and inhibitory couplings. A key feature of this structure is that the eigenvectors of $J$ are non-orthogonal and tend to become nearly parallel under the balance condition $\zeta\equiv\frac{\gamma_{\mu}-\gamma_{\nu}}{\gamma_{\mu}+\gamma_{\nu}}\approx 0$. In this regime, the eigenbasis poorly spans the full phase space: directions orthogonal to the dominant eigenvectors are underrepresented, and any perturbation with a component in these directions requires large contributions from the available (nearly aligned) eigenvectors. As a result, even small imbalances along $\Delta$ can lead to disproportionately large transient excursions in $\Sigma$, despite the presence of negative eigenvalues and overall linear stability. This reflects the hallmark of non-normal dynamics: strong directional sensitivity and transient amplification that escape standard eigenvalue-based stability analysis.
@@
$\frac{dx(t)}{dt}$ $=-\alpha x+(1-xf(s)+\sqrt{\alpha x+(1-x)f(s)}\eta_{x}(t)$ (3)
@@
where $\eta_{x,y}$ are uncorrelated Gaussian white noises, with amplitude $\sigma$ that depends on the network-size $\sigma\propto 1/\sqrt{N}$ and some time dependencies have been omitted for simplicity in the notation.
@@
Structural non-normal effects: feedforward and feedback network motifs.
To investigate how non-normal dynamics generalize from individual excitatory-inhibitory (x/y) units to interconnected circuits, we now analyze simple network motifs composed of multiple coupled x/y populations. Our goal is to understand how the non-normal effects identified at the single-unit level are shaped by the architecture of inter-unit interactions.
We focus on a minimal yet informative motif comprising three local x/y populations, as illustrated in Fig.1. Inter-population connectivity is mediated exclusively by excitatory neurons: these not only project to excitatory neurons in other units but also to inhibitory neurons across populations. In contrast, inhibitory neurons act only locally within each population, consistent with biological observations [26]. Because inter-population coupling is purely excitatory, we summarize the global architecture using the $\times 3$ 3 adjacency matrix $A$, which captures the excitatory-to-excitatory connectivity pattern:
@@
The parameter $0\leq c\leq\epsilon$ interpolates between a symmetric cycle ($c=\epsilon$) and a purely feedforward structure ($c=0$), thus controlling the degree of non-normality. The full $\times 6$ 6 adjacency matrix, which includes both excitatory and inhibitory neurons, is not shown explicitly. However, it consists of this $\times 3$ 3 excitatory submatrix duplicated in the excitatory-to-inhibitory blocks, while the inhibitory subnetwork remains purely diagonal, since inhibition is strictly local.
Figure 1: Inter-populations interactions. a) Schematic representation of the generic non-reciprocal motif. X and Y respectively stand for the fraction of active excitatory and inhibitory populations. Excitatory couplings are represented in dark red, while inhibitory interactions are shown in dark green. Interactions between populations are illustrated by light green arrows. Two of these inter-population connections are modulated by the parameter $\epsilon$, while only one is modulated by the parameter $0\leq c\leq\epsilon$, according to Eq. 4. This allows us to control the behavior of the system between the homogeneous loop and the feed-forward case. b) Schematic representation of the model following from the transformation ($\Sigma$ = x+y, $\Delta$ = x-y). The inter-populations interactions are drawn in orange for the cyclic setup and in red for the feedforward one.
The full set of coupled Langevin equations that we consider is then:
@@
where $\gamma_{\ell}$ sets the excitatory coupling strength between units. To fix the ideas we will consider $f(s)=\Theta(s)\tanh{s}$, where $\Theta(s)$ is the Heaviside function, or its derivable counterpart $f_{\xi}(s)=\frac{1}{2}\left(1+\frac{2}{\pi}\arctan\left(\frac{s}{\xi}\right)\right)\tanh{s}$, with for $\xi<<1$).
We focus primarily on motifs composed of three x/y populations, which already capture rich and nontrivial dynamical behavior. Generalization to $n$ coupled populations is presented in the Supplemental Material (SM).
The Jacobian of this set of equations develops complex eigenvalues when three or more coupled units are considered, as detailed in Sec.III.2, allowing us one to account for more interesting transient and time-dependent behaviors such as noise-induced oscillations [27]. The paper is organized as follows: in Section III.1 we analyze the steady state of Eq. (5) for the cyclic and feedforward motifs; in Section III.2 we perform linear analysis to gain insight on the features of the orbits around the stable fixed point; in Section III.3 we examine the reactivity properties and in Section 4 we relate them to the behavior of the stochastic system.
@@ III.1 Steady state analysis
First, we analyze the type of bifurcation that the deterministic system undergoes near the zero fixed point, i.e. when it shifts from inhibition-dominated to excitation-dominated, focusing on the comparison between the two limit architectures introduced above. Then, we study the controllability of the system, by quantifying how rapidly the network escapes from the inactive phase using the excitatory coupling $\gamma_{\mu}$ as a control parameter.
Figure 2: Bifurcation diagram and scaling of the fixed points. Panel a) shows the bifurcation diagram for basal population $x_{1}$ obtained via Runge-Kutta integration. The parameters $\alpha=0.1$, $\gamma_{\nu}=0.35$ and $\gamma_{\ell}=0.2$ have been used for several setups. The orange line confirms the bifurcation point at $\gamma_{\mu}=\gamma_{\mu}^{c,\epsilon}=0.25$ for the cyclic system. The decoupled diagram, with bifurcation point predicted at $\gamma_{\mu}=\gamma_{\mu}^{c}=0.45$, is shifted to make $\gamma_{\mu}^{c,\epsilon}=\gamma_{\mu}^{c}$. Remarkably, the cyclic and the decoupled lines coincide over the whole curve. Eventually, several values of $c$ are evaluated, up to $c=0$ (feed-forward setup), with expected bifurcation point $\gamma_{\mu}^{c,0}=0.45$. In the other panels, the scaling of the fixed point around the bifurcation. The behavior for $c=\epsilon/2$ in panels c), d) and e) shows how the linear scaling extends beyond $c=\epsilon$ with reduced but reliable approximation for each population of the cyclic setup. Panel b) refers to the feed-forward case.
For the sake of simplicity, let us start by discussing the decoupled setup ($\gamma_{\ell}=0$), studied in [13, 7]. In this case, the units are trivially just copies of the same system, i.e. $x_{i}=x_{j},y_{i}=y_{j}$ for all couples of indices $i,j$. More interestingly, the equilibrium points satisfy the identities: $\bar{x}_{i}=\bar{y}_{i}$, as anticipated above. We now would like to extend the validity of these equalities for a certain region around the fixed points. As we will discuss in detail in Section III.2, the system is strongly non-normal and both eigenvectors of the Jacobian align along the bisector of the $x_{i}-y_{i}$ plane. This allows us to suppose that for $t>>\infty$, $x_{i}(t)\sim y_{i}(t)$. Therefore, the dynamical system reduces to a one-dimensional single equation.
[…]
@@
which is the normal form for a transcritical bifurcation [28]. Then, the non-zero solution of this equation is
@@
and the bifurcation occurs at $\gamma_{\mu}^{c}=\gamma_{\nu}+\alpha$. Then, linearizing around this bifurcation point we get the scaling $\bar{\text{x}}\sim\frac{1}{\alpha}\left(\gamma_{\mu}-\gamma_{\mu}^{c}\right)$.
Cyclic coupling

### v2

@@ Abstract
Non-reciprocal interactions are widespread in neural circuits, arising both from asymmetric local excitation–inhibition dynamics and from directed coupling between distinct populations. Here we investigate how these two sources of asymmetry jointly shape network activity using a modified Wilson-Cowan framework in which each node is an excitatory/inhibitory population operating in a strongly non-normal regime and nodes are linked by directed excitatory projections. We analyze minimal three-node motifs that interpolate between a feedback (cyclic) architecture and a purely feedforward chain. By combining steady-state and bifurcation analysis with linear stability and reactivity measures, we show that feedback coupling largely maps onto an effective parameter shift relative to the uncoupled case, whereas feedforward coupling produces qualitatively different activation properties, including distinct scaling across stages and a genuine enhancement of reactive transient amplification. Incorporating intrinsic finite-size noise, we find fluctuation-driven switching between inactive and active states and motif-dependent spatiotemporal organization: feedback promotes coherent, correlated activation across nodes, while feedforward networks display sequential, step-like propagation. Together, these results provide a minimal framework for understanding how multi-scale non-reciprocities control transient responses and noise-driven dynamics in coupled neural populations.
@@ I Introduction
In this paper, we distinguish two manifestations of physical non-reciprocity in complex systems—arising either from asymmetric local excitatory-inhibitory dynamics at a local level or from directed network architecture—which often co-occur but are conceptually separate. We use non-normal linear operators as a convenient mathematical framework to describe both types of non-reciprocity and analyze their dynamical consequences. Recall that a matrix is non-normal when it does not commute with its transpose (for real matrices) or adjoint (for complex matrices), e.g. $AA^{\top}\neq A^{\top}A$. In such cases, its (right) eigenmodes are generally non-orthogonal, and this non-orthogonality can enable strong transient amplification even when all eigenvalues indicate asymptotic stability.
@@
The first form we address is local dynamical non-normality, which arises at the level of the internal dynamics of individual units or subsystems. For example, consider an island with a predator population and a prey population. The local Jacobian (stability matrix) around a stationary point of the dynamics can be non-normal, so that even if the fixed point is asymptotically stable, the system may exhibit large transient responses to small perturbations—an effect that can be enhanced by continual disturbances or noise [30, 29, 35, 42, 37, 36]. Similarly, in neural circuits, interacting excitatory and inhibitory neuronal populations often operate in regimes where the effective linearized dynamics is strongly non-normal, allowing transient amplification of input fluctuations and supporting fast signal processing, selectivity, and flexible population coding [19, 2, 3, 1, 27, 12, 6, 39, 4].
@@
The second form is what we term structural non-normality, which stems from network-level asymmetry in how units influence one another at a larger scale. In many real-world systems, especially biological, and engineered networks, the connectivity that mediates inter-unit interactions is directed and typically not symmetric. Ecological and social systems, for example, often exhibit trophic or influence hierarchies, in which some quantity flows preferentially along directed pathways [20, 23]. Neural circuits likewise are not fully bidirectional: higher-order cortical areas send feedback to early sensory regions, but not necessarily with equal strength in the opposite direction [17, 14, 32, 18]. Such directed structure can (but need not) generate a non-normal connectivity operator; when it does, it can promote directional amplification, long-range correlations, wave-like responses, and flow-like propagation of activity across the network.
Here, we test the hypothesis that embedding locally non-normal units within globally non-reciprocal architectures yields collective behaviors stronger or distinct from those produced by either source of asymmetry alone. Specifically, we explore how the interplay between local dynamical non-normality (within E-I populations) and global structural non-normality (across network tiers) determines the reactivity of the system. We show that while feedback architectures largely rescale the effective parameters of local units, feedforward motifs fundamentally alter the scaling laws of the system, introducing a hierarchy of timescales and amplifying transient responses. In other words, when locally asymmetric (and potentially non-normal) subsystems (e.g., excitatory/inhibitory microcircuits) are embedded in a directed large-scale architecture (e.g., feedforward/feedback pathways), the resulting macroscopic dynamics may exhibit emergent phenomena not captured by standard linear-response intuition for symmetric interactions. These may include enhanced stochastic amplification, novel transient modes, and selective propagation of fluctuations and signals across the network.
@@
As a test of this hypothesis, here we analyze a minimal yet general neural model consisting of multiple interacting dynamical units, each described by a non-normal local E-I system, coupled through a network, that in general is non-reciprocal. Using tools from linear stability analysis, stochastic dynamics, and spectral theory, we explore how the composition of non-reciprocities —dynamical and structural— governs the system’s susceptibility to noise, its capacity to transmit signals, and its repertoire of emergent dynamical states.
@@
While our focus is primarily on neural systems —where both forms of non-reciprocity are particularly notorious— the conceptual framework we propose is broadly applicable to other domains, including active materials, biological networks, and collective decision-making systems. In all these settings, non-reciprocity functions as a structural and dynamical source of complexity, and elucidating its distinct manifestations is crucial for understanding the behavior of high-dimensional, complex systems.
@@ II Wilson-Cowan model for neural populations
II.1 Dynamical non-normal effects in E/I populations.
@@
$\begin{aligned} \frac{dx(t)}{dt}&=-\alpha x(t)+(1-x(t))f(s)\\ \frac{dy(t)}{dt}&=-\alpha y(t)+(1-y(t))f(s),\end{aligned}$ (1)
where $x(t)$ and $y(t)$ represent the levels of activation of excitatory and inhibitory neuronal population, respectively. The parameter $\alpha$ represents the rate of spontaneous activity decay, while $s$ denotes the net incoming input, $s(t)=\gamma_{\mu}x(t)-\gamma_{\nu}y(t)+h$, which is simply the sum of all synaptic excitatory and inhibitory inputs, weighted by their respective synaptic strengths plus a constant external input, $h$. To reduce the dimensionality of the parameter space, all synaptic weights of a given type have been set to the same value: $(\gamma_{\mu}$ for excitation and $\gamma_{\nu}$ for inhibition). Finally, $f(s)$ is a sigmoid non-negative response function. The autonomous system ($h=0$) displays two distinct dynamical regimes: a quiescent (inactive) phase, where the mean activity vanishes, $\bar{x}=\bar{y}=0$, and an active phase characterized by nonzero steady-state activity $\bar{x}=\bar{y}>0$. These phases are separated by a bifurcation that occurs at the critical point $\gamma_{\mu}=\gamma_{\nu}+\alpha$: below this threshold, the system remains in the quiescent state, while above it, activity becomes self-sustained (see below).
It is useful to make the change of variables $\Sigma=(x+y)/2$ and $\Delta=(x-y)/2$, to consider the overall activation and the differential one between excitation and inhibition, respectively. Linearizing around the active-state fixed point $(\bar{\Sigma},\bar{\Delta}\neq 0)$, one obtains a strictly upper-triangular Jacobian matrix, which is manifestly non-normal:
@@
where the eigenvalues are given by $\lambda_{1}=\alpha+f(\bar{s})+(1-\bar{\Sigma})\omega_{0}f^{\prime}(\bar{s})$ and $\lambda_{2}=\alpha+f(\bar{s})$, with $\bar{s}=\omega_{0}\bar{\Sigma}+h$. The feedforward term $\omega_{\text{ff}}=(1-\bar{\Sigma})(\gamma_{\mu}+\gamma_{\nu})f^{\prime}(\bar{s})$ governs the asymmetric influence of $\Sigma$ on $\Delta$, while the parameter $\omega_{0}=\gamma_{\mu}-\gamma_{\nu}$ quantifies the difference between excitatory and inhibitory couplings. A key feature of this structure is that the eigenvectors of $J$ are non-orthogonal [5, 12] and tend to become nearly parallel under the condition $\zeta\equiv\frac{|\gamma_{\mu}-\gamma_{\nu}|}{\gamma_{\mu}+\gamma_{\nu}}\approx 0$, which describes balance between excitation and inhibition. In this regime, the eigenbasis poorly spans the full phase space: directions orthogonal to the dominant eigenvectors are poorly represented, and any perturbation with a component in these directions requires large contributions from the available (nearly aligned) eigenvectors. As a result, even small fluctuations along $\Delta$ can lead to giant transient excursions in $\Sigma$, despite the presence of negative eigenvalues and overall linear stability [5, 12, 8]. This reflects the hallmark of non-normal dynamics: strong directional sensitivity and transient amplification that escape standard eigenvalue-based stability analysis.
@@
$\frac{dx(t)}{dt}$ $=-\alpha x+(1-x)f(s)+\sqrt{\alpha x+(1-x)f(s)}\eta_{x}(t)$ (3)
@@
where $\eta_{x,y}$ are uncorrelated Gaussian white noises, with amplitude $\sigma$ that depends on the network-size $\sigma\propto 1/\sqrt{N}$ (note that some time dependencies have been omitted for simplicity in the notation).
Figure 1: Inter-population interactions. (a): Schematic representation of the generic non-reciprocal motif. $X$ and $Y$ denote the fractions of active excitatory and inhibitory populations, respectively. Local excitatory couplings are shown in dark green, while inhibitory interactions are indicated in dark red. Light green arrows illustrate inter-population interactions. Two of these connections (represented in light blue) are modulated by the parameter $c$, with $0\leq c\leq\epsilon$, as defined in Eq. (4). This modulation allows to interpolate between the homogeneous loop and the feedforward configuration. (b): Schematic representation of the model in the transformed variables ($\Sigma=x+y$, $\Delta=x-y$). In the transformed variables (b) the feedforward configuration includes both within-unit (orange) and between-units (red) couplings, while the cyclic network only includes within-unit (orange) interactions, i.e. the three units are disjoint.
@@
II.2 Structural non-normal effects: feedforward and feedback network motifs.
To investigate how non-normal dynamics extend from individual excitatory-inhibitory (x/y) units to interconnected circuits, we analyze simple network motifs composed of multiple coupled x/y populations. Our aim is to understand how the non-normal features identified at the single-unit level are modulated by the architecture of inter-unit interactions.
Here we focus on motifs composed of three x/y populations (or "units" or "areas"), schematically illustrated in Fig. 1, which already capture rich and nontrivial dynamical behavior (generalization to $n$ coupled populations and more complex motifs are analyzed in the Supplemental Material). Inter-population connectivity is mediated exclusively by excitatory neurons projecting to both excitatory and inhibitory neurons in other populations. In contrast, inhibitory neurons act only locally within each population, in accordance with biological observations [21]. Because inter-population coupling is purely excitatory, the global architecture can be compactly described by a $3\times 3$ adjacency matrix $A$, which encodes the excitatory-to-excitatory connectivity pattern:
@@
The parameter $0\leq c\leq\epsilon$ interpolates between a symmetric cycle ($c=\epsilon$) and a purely feedforward structure ($c=0$), thereby controlling the degree of non-normality. The full $6\times 6$ adjacency matrix, which includes both excitatory and inhibitory neurons, is not shown explicitly (but note that it is constructed from the previous $3\times 3$ excitatory submatrix: this submatrix is duplicated in the excitatory-to-inhibitory blocks, while the inhibitory subnetwork remains purely diagonal, reflecting the strictly local nature of inhibition).
The full set of coupled Wilson-Cowan equations that we consider is then:
@@
with $i=,1,2$ or $3$ and where $\gamma_{\ell}$ sets the excitatory coupling strength between units, and $f(s)=\Theta(s)\tanh{s}$, with $\Theta(s)$ the Heaviside function.
The Jacobian of this system develops complex eigenvalues when three or more coupled units are considered, as detailed in Sec. III.2, allowing us to account for interesting transient and time-dependent behaviors, such as noise-induced oscillations [13]. These $\epsilon$-dependent minimal motifs allow us to isolate how architectural directionality modifies bifurcation structure, eigenvector geometry, reactivity, and noise-driven dynamics, while keeping the local non-normal mechanism fixed. The rest of the paper is organized as follows: in Sec. III.1 we analyze the steady state of Eq. (5) for the cyclic and feedforward motifs; in Sec. III.2 we perform a linear stability analysis to characterize the features of the orbits around the stable fixed point; in Sec. III.3 we examine the reactivity properties; and in Sec. III.4 we relate these properties to the behavior of the stochastic system.
@@ III.1 Steady state analysis
First, we analyze the type of bifurcation that the deterministic system undergoes near the vanishing-activity fixed point, i.e., when the system transitions from inhibition-dominated to excitation-dominated dynamics, focusing on the comparison between the two limit architectures introduced above. Subsequently, we study the controllability of the system by quantifying how abruptly the network escapes from the inactive phase, using the excitatory coupling $\gamma_{\mu}$ as a control parameter.
Figure 2: Bifurcation diagram and scaling of the fixed points. Panels (a–c) Scaling of the fixed points around the bifurcation for $x_{1},x_{2}$, and $x_{3}$ respectively. The scaling is the same for the decoupled case (purple), for $c=\epsilon/2$ (light blue) and $c=\epsilon$ (orange). (d) Scaling of the fixed points around the bifurcation for the Feedforward case ($c=0$). Note the change of the slope as expected by theoretical results, depending on the population (see legend). (e) Bifurcation diagram for the basal population $x_{1}$, obtained via Runge-Kutta integration. Parameters $\alpha=0.1$, $\gamma_{\nu}=0.35$, and $\gamma_{\ell}=0.2$ are used across several setups. The orange line marks the bifurcation point at $\gamma_{\mu}=\gamma_{\mu}^{c,\epsilon}=0.25$ for the cyclic system. The decoupled diagram, with bifurcation point predicted at $\gamma_{\mu}=\gamma_{\mu}^{c}=0.45$, is shifted so that $\gamma_{\mu}^{c,\epsilon}=\gamma_{\mu}^{c}$, resulting in a near-perfect overlap of the cyclic and decoupled curves. Several values of $c$ are also evaluated, up to $c=0$ (feedforward setup), with the expected bifurcation point $\gamma_{\mu}^{c,0}=0.45$.
For the sake of simplicity, we begin by discussing the decoupled setup ($\gamma_{\ell}=0$), previously studied in [5, 12]. In this case, the units are trivially copies of the same system, i.e., $x_{i}=x_{j}$ and $y_{i}=y_{j}$ for all pairs of indices $i,j$ and the equilibrium points satisfy $\bar{x}_{i}=\bar{y}_{i}$. We now aim to extend the validity of these equalities to a region surrounding the fixed points. As discussed in detail in Sec. III.2, the system is strongly non-normal, and both eigenvectors of the Jacobian align along the bisector of the $x_{i}$$y_{i}$ plane. This alignment allows us to assume that, for $t\gg 1$, $x_{i}(t)\sim y_{i}(t)$. Consequently, the dynamical system effectively reduces to a one-dimensional equation. In particular, the normal form around the bifurcation point $\text{x}=0$, with $h=0$, can be written as
@@
which formally corresponds to the normal form for a transcritical bifurcation [34]. The non-zero solution of this equation is
@@
and the bifurcation occurs at the critical value $\gamma_{\mu}^{c}=\gamma_{\nu}+\alpha$. Then, linearizing around this bifurcation point one gets $\bar{\text{x}}\sim\frac{1}{\alpha}\left(\gamma_{\mu}-\gamma_{\mu}^{c}\right)$.
III.1.1 Cyclic coupling